# DPOT: Auto-Regressive Denoising Operator Transformer for Large-Scale PDE Pre-Training

### v2

@@ Abstract
Pre-training has been investigated to improve the efficiency and performance of training neural operators in data-scarce settings. However, it is largely in its infancy due to the inherent complexity and diversity, such as long trajectories, multiple scales and varying dimensions of partial differential equations (PDEs) data. In this paper, we present a new auto-regressive denoising pre-training strategy, which allows for more stable and efficient pre-training on PDE data and generalizes to various downstream tasks. Moreover, by designing a flexible and scalable model architecture based on Fourier attention, we can easily scale up the model for large-scale pre-training. We train our PDE foundation model with up to 0.5B parameters on 10+ PDE datasets with more than 100k trajectories. Extensive experiments show that we achieve SOTA on these benchmarks and validate the strong generalizability of our model to significantly enhance performance on diverse downstream PDE tasks like 3D data. Code is available at https://github.com/thu-ml/DPOT.
@@ 2.1 Neural Operators
Further progress has been made through new architectures utilizing transformers (Cao, 2021; Li et al., 2022a; Hao et al., 2023; Wu et al., 2023; Ovadia et al., 2023), integrating techniques like patchification and linear attention mechanism. For instance, the GK-Transformer (Cao, 2021), OFormer (Li et al., 2022a), and GNOT (Hao et al., 2023) have shown promise on problems with irregular geometries. AFNO (Guibas et al., 2021) proposes efficient attention by applying FFT to the data and mixes features in the frequency domain inspired by FNO, which achieves a remarkable capacity with a low memory and computational cost. It is then modified and deployed for large-scale climate forecasting (Pathak et al., 2022). Despite these significant efforts, they need to be individually trained for specific tasks and require much domain-specific data, highlighting the need to improve data efficiency in this field.
@@ 3.5 Analysis of Our Model
Theoretical Analysis. We present a theoretical analysis to show the expressivity of our model architecture. Specifically, we prove the following universal approximation theorem for Fourier attention layers, and the proof is deferred to Appendix C.2.
@@ 4 Experiments
Training and Evaluation. For our training process, we selected $T=10$ timesteps to predict the next frame, aligning with the original settings of most datasets.
The model sizes are specified in Appendix A.1.
For models of all scales, we utilized the AdamW optimizer with a learning rate of $1\times 10^{-3}$ and trained the models for 1000 epochs. Our model was trained on servers equipped with eight A800 GPUs, each with 80 GB memory. Unless otherwise stated, we use $l_{2}$ relative error (L2RE) to measure the quality of the prediction, which is commonly used and can be referred to in previous works (Li et al., 2020).
L2RE Params FNO-$\nu$ PDEBench CNS,DR,SWE PDEArena CFDBench
@@ 4.3 Main Results
Table 1 presents the results of our main experiments, i.e., the L2RE on the test sets of datasets in pre-training.
The first part of the table shows the results of pre-trained and non-pre-trained small models ($\leqslant 30$M). We see that our model successfully learns from multiple PDE datasets simultaneously using pre-training. We achieve state-of-the-art performance on 9 out of 12 datasets with an average improvement up to 52% ($0.116$ to $0.0553$ on FNO-$1e-4$). This indicates that our auto-regressive denoising pre-training strategy is effective for highly complex and heterogeneous PDE data. Specifically, our model outperforms most datasets with roughly equivalent model parameters. Compared to MPP, our model supports more datasets, i.e., three FNO datasets, two PDEArena datasets, and CFDBench dataset, and performs the best on most datasets except for two small subsets, i.e., PDB-SWE and PDB-DR. This demonstrates that given similar model sizes, the expressive power of our model’s structure is more powerful in operator learning.
@@
The last part of the table displays the performance of differently sized pre-trained DPOTs after fine-tuning. The “-200” and “-500” suffixes denote fine-tuning on each subset for 200 and 500 epochs, respectively. We see that larger models also yield better fine-tuning results. The pre-training plus fine-tuning of DPOT achieves SOTA across all results, while the SOTA of 10 out of 12 tasks is achieved by fine-tuning the largest DPOT model (0.5B). It reduces L2RE for more than 50% on 8 out of 12 tasks compared with all zero-shot methods. This suggests that larger pre-trained models extract more knowledge from the datasets and provide significantly better initialization for fine-tuning. Moreover, we also observe that fine-tuning more steps is usually better than fewer steps, which shows there exists a trade-off between the computational cost and performance.
@@ 4.4 Downstream Tasks Experiments
$*{}^{*}$ Turbulence 3D dataset $*{}^{*}$
@@
Table 2: Experimental results of fine-tuning on downstream tasks. The first row shows the challenges of the downstream tasks. For steady-state PDE with $*{}^{*}$, we adopt relative median $L^{1}$ error (L1-Median) following the original work (Raonić et al., 2023). Lower error means better performance, and we bold the best results.
@@
To validate the effectiveness of our pre-trained model as a foundation for enhancing various types and complexities of PDE downstream tasks, we conducted the following experiments to explore the broader usage of the model. We transferred the weights of the Fourier attention block from our pre-trained model to the models for downstream tasks as initial values. To demonstrate that our model learns knowledge that generalizes to diverse PDEs, we choose our downstream tasks, including trajectory prediction for 2D high-resolution turbulence flow, 3D Navier-Stokes equations, and operator learning for steady-state PDEs. The results are displayed in Table 2. The DPOT-Vanilla denotes the results of training the DPOT from scratch for 500 epochs, while DPOT-FT represents training for the same number of steps but using pre-trained weights for initialization. We have the following observations from the results.
@@ 4.6 Ablation Experiments
Here, we conduct ablation studies to investigate the influence of some hyperparameters and modules in our method. In these experiments, we train small-scale models on PDEBench datasets and compare the averaged zero-shot performance on corresponding test datasets.
@@
Influence of noise injection level. Here, we discuss the impact of the magnitude of noise injection on network training and generalization. As mentioned before, we observed that introducing noise into the input during auto-regressive training enhances the model’s robustness and stability for long trajectories in the test set. We experiment with noise levels $\varepsilon\in\{0,5\times 10^{-5},5\times 10^{-4},5\times 10^{-3},5\times 10^{-2}\}$, and the results are shown in Table 4. We find that the training error increases with the intensity of the injected noise, indicating that data with added noise becomes more challenging to learn. However, the test error shows a tendency to decrease from $0.0178$ to $0.0152$ before increasing with the addition of noise. We could view the noise injection as a regularization, and an appropriate level of noise benefits the generalization.
@@ 5 Conclusion
In this paper, we proposed the DPOT for large-scale PDE pre-training. We designed the new auto-regressive denoising strategy and a new model architecture based on Fourier attention. We pre-trained a model up to 0.5B with 10+ PDE datasets. Then, we conducted extensive experiments on multiple PDE benchmarks and downstream PDE tasks. Our work is a pioneering work in large-scale exploration of the scaling behavior of PDE pre-training. It shows great potential for future applications as downstream tasks in practical problems.
@@ A.1 Model sizes and training details
Pre-training. We selected four models of varying sizes, i.e., DPOT-Tiny, DPOT-Small, DPOT-Medium, and DPOT-Large. The specific parameters of these models are shown in the table below. For the pre-training stage, we set the learning rate to $1\times 10^{-3}$ and used a One-cycle learning rate schedule over 1000 epochs, with the first 200 epochs as the warm-up phase. The AdamW optimizer was employed with a weight decay of $1\times 10^{-6}$ and momentum parameters $(\beta_{1},\beta_{2})=(0.9,0.9)$. Training was conducted using eight A800 GPUs, with a total batch size of 160. The patch size was set to 8. For datasets with long trajectories, such as PDB-DR and PDB-SWE, we assigned a weight of $w=3$, while for others, the weight was set to 1.
@@ Appendix B Supplementary Experimental Results
B.1 Timing experiments
@@
B.2 Comparison with AFNO architecture
@@
B.3 Performance under varying resolutions
@@ Appendix C Proof of Universal Approximation of Fourier Attention
In the following section, we will demonstrate the Universal Approximation capability of DPOT. This means that the model possesses ample expressive ability, capable of fitting a large class of operators and adapting to the required precision. Specifically, we show the core layers, i.e. Fourier attention layers have the capacity to approximate any continuous operators. Then similar to (Kovachki et al., 2021), we could extend results to discretized versions with patchfication layers.
@@ C.1 Preliminaries and Definition
To begin with, we will now present the essential definitions and notations required for the proof.
@@
As DPOT mainly focuses on PDE problems defined on a rectangular domain, we will concentrate on functions defined on periodic torus $\mathbb{T}^{d}=[0,2\pi]^{d}$. Given the domain, we can define the Fourier transform as:
$\mathcal{F}(v)(k)=\frac{1}{(2\pi)^{d}}\int_{\mathbb{T}^{d}}v(x)e^{-i\langle k,x\rangle}\text{d}x,\ \ \forall k\in\mathbb{Z}^{d}$ (22)
for any function $v\in H^{s}(\mathbb{T}^{d})$. We also denote $\hat{v}=\mathcal{F}(v)$ and $\hat{v}_{k}=\mathcal{F}(v)(k)$ for the $k$-th Fourier coefficient.
Given $\hat{v}$, we can also define inverse discrete fourier transform as:
$\mathcal{F}^{-1}(\hat{v})(x)=\sum_{k\in\mathbb{Z}^{d}}\hat{v}_{k}e^{i\langle k,x\rangle},\forall x\in\mathbb{T}^{d}$ (23)
In practical terms, computing and storing all the Fourier coefficients for a given function is unfeasible due to their infinite length. Therefore, a common practice is to truncate the components with frequencies exceeding N. Formally, we can define Fourier wavenumbers as:
$\mathcal{K}_{N}=\{k\in\mathbb{Z}^{d}\colon|k|_{\infty}\leq N\},$ (24)
and define a projection operator to truncate the high frequencies part of a function as:
$\mathcal{P}_{N}(v)(x)=\mathcal{F}^{-1}(\mathcal{F}(v)(k)\cdot\mathbf{1}_{\mathcal{K}_{N}}(k))$ (25)
where $1_{\mathcal{K}_{N}}(k)$ takes $1$ when $k\in\mathcal{K}_{N}$ and 00 otherwise.
@@
Now, with these preparations in place, we can formally define the DPOT. A DPOT is a neural operator $\mathcal{N}$ defined between input function space $H^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{in}})$ and output space $H^{s^{\prime}}(\mathbb{T}^{d};\mathbb{R}^{d_{out}})$. Here, $H^{s}$ represents the Sobolev space of order $s$, and $d_{in},d_{out}$ are the number of channels of the input and output functions. The structure of DPOT can be described as:
$\mathcal{N}=\mathcal{Q}\circ\mathcal{L}_{L}\circ\mathcal{L}_{L-1}\circ\cdots\circ\mathcal{L}_{1}\circ\mathcal{R},$ (26)
where $\mathcal{R}$ is the lifting operator mapping the codomain of input function into hidden space $\mathbb{R}^{d_{hid}}$, which is actually the composition of Embedding layer and Temporal Aggregation Layer in the context of DPOT:
$\mathcal{R}(v)(x)=R(v(x)),\quad R\in C_{\infty}(\mathbb{R}^{d_{in}};\mathbb{R}^{d_{hid}})$ (27)
and $\mathcal{Q}$ is the projection operator, mapping hidden space into output space $\mathbb{R}^{d_{out}}$
$\mathcal{Q}(v)(x)=Q(v(x)),Q\in C_{\infty}(\mathbb{R}^{d_{hid}};\mathbb{R}^{d_{out}})$ (28)
@@
Finally, the Fourier attention layers $\mathcal{L}_{l}$ of DPOT can be defined as:
$\mathcal{L}_{l}(v)(x)=v(x)+\mathcal{M}_{l}\circ\mathcal{F}^{-1}(K_{l}(k)(\mathcal{F}(\mathcal{P}_{N}v)(k)))(x),$ (29)
where $\mathcal{M}_{l}(v)(x)=M_{l}(v(x))$ is a 2-layer MLP with GELU activation $\sigma$, applying pointwise transform to the input, and $K_{l}:\mathbb{Z}^{d}\times\mathbb{C}^{d_{hid}}\to\mathbb{C}^{d_{hid}}$ is a 2-layer Multi-Head Frequency Feature MLP, defines a non-linear transform on the Fourier coefficients.

### v4

@@ Abstract
Pre-training has been investigated to improve the efficiency and performance of training neural operators in data-scarce settings. However, it is largely in its infancy due to the inherent complexity and diversity, such as long trajectories, multiple scales and varying dimensions of partial differential equations (PDEs) data. In this paper, we present a new auto-regressive denoising pre-training strategy, which allows for more stable and efficient pre-training on PDE data and generalizes to various downstream tasks. Moreover, by designing a flexible and scalable model architecture based on Fourier attention, we can easily scale up the model for large-scale pre-training.We train our PDE foundation model with up to 1B parameters on 10+ PDE datasets with more than 100k trajectories. Extensive experiments show that we achieve SOTA on these benchmarks and validate the strong generalizability of our model to significantly enhance performance on diverse downstream PDE tasks like 3D data. Code is available at https://github.com/thu-ml/DPOT.
@@ 2.1 Neural Operators
Further progress has been made through new architectures utilizing transformers (Cao, 2021; Li et al., 2022a; Hao et al., 2023; Wu et al., 2023; Ovadia et al., 2023), integrating techniques like patchification and linear attention mechanism. For instance, the GK-Transformer (Cao, 2021), OFormer (Li et al., 2022a), and GNOT (Hao et al., 2023) have shown promise on problems with irregular geometries. AFNO (Guibas et al., 2021) proposes efficient attention by applying FFT to the data and mixes features in the frequency domain inspired by FNO, which achieves low memory and computational cost similar to MLP-Mixer(Tolstikhin et al., 2021). It is then modified and deployed for large-scale climate forecasting (Pathak et al., 2022). Despite these significant efforts, they need to be individually trained for specific tasks and require much domain-specific data, highlighting the need to improve data efficiency in this field.
@@ 3.5 Analysis of Our Model
Theoretical Analysis. We present a theoretical analysis to show the expressivity of our model architecture. Specifically, we prove the following universal approximation theorem for Fourier attention layers, and the proof is deferred to Appendix LABEL:proof-thm.
@@ 4 Experiments
Training and Evaluation. The model sizes are specified in Appendix A.1. For models of all scales, we utilized the AdamW optimizer with a learning rate of $1\times 10^{-3}$ and trained the models for 1000 epochs. Our model was trained on servers equipped with eight A800 GPUs, each with 80 GB memory. Unless otherwise stated, we use $l_{2}$ relative error (L2RE) to measure the quality of the prediction, which is commonly used and can be referred to (Li et al., 2020).
L2RE Params FNO-$\nu$ PDEBench CNS-$(\eta,\zeta)$,DR,SWE PDEArena CFDBench
@@ 4.3 Main Results
Table 1 presents the results of our main experiments, i.e., the L2RE on the test sets of datasets in pre-training. The parameter in the second row means the parameter for the PDE dataset, e.g. $51e-$ 5 means the viscosity $\nu=1e-5$ in FNO dataset (Li et al., 2020) and $1,0.1$ means $(\eta,\zeta)=(1,0.1)$ in PDEBench (Takamoto et al., 2022).
The first part of the table shows the results of pre-trained and non-pre-trained small models ($\leqslant 30$M). We see that our model successfully learns from multiple PDE datasets simultaneously using pre-training. We achieve state-of-the-art performance on 9 out of 12 datasets with an average improvement up to 52% ($0.116$ to $0.0553$ on FNO-$1e-4$). This indicates that our auto-regressive denoising pre-training strategy is effective for highly complex and heterogeneous PDE data. Specifically, our model outperforms most datasets with roughly equivalent model parameters. Compared to MPP, our model supports more datasets and performs the best on most datasets except for two small subsets, i.e., PDB-SWE and PDB-DR. This demonstrates that given similar model sizes, the expressive power of our model’s structure is more powerful in operator learning.
@@
The last part of the table displays the performance of differently sized pre-trained DPOTs after fine-tuning. The “-200” and “-500” suffixes denote fine-tuning on each subset for 200 and 500 epochs, respectively. We see that larger models also yield better fine-tuning results. The pre-training plus fine-tuning of DPOT-L(0.5B) achieves SOTA in 9 out of 12 tasks. It reduces L2RE for more than 50% on 8 out of 12 tasks compared with all zero-shot methods. This suggests that larger pre-trained models extract more knowledge from the datasets and provide significantly better initialization for fine-tuning. Moreover, we also observe that fine-tuning more steps is usually better than fewer steps, which shows there exists a trade-off between the computational cost and performance.
@@ 4.4 Downstream Tasks Experiments
L2RE/L1-Medium Turbulence 3D dataset Steady
@@
Table 2: Experimental results of fine-tuning on downstream tasks. The first row shows the challenges of the downstream tasks. For steady-state PDE with , we adopt relative median $L^{1}$ error (L1-Median) following the original work (Raonić et al., 2023). Lower error means better performance, and we bold the best results.
@@
To validate the effectiveness of our pre-trained model as a foundation for enhancing various types and complexities of PDE downstream tasks, we conducted the following experiments to explore the broader usage of the model. We transferred the weights of the Fourier attention block from our pre-trained model to the models for downstream tasks as initial values. We choose our downstream tasks, including trajectory prediction for 2D high-resolution turbulence flow, 3D Navier-Stokes equations, and operator learning for steady-state PDEs. The results are displayed in Table 2. The DPOT-Vanilla denotes the results of training the DPOT from scratch for 500 epochs, while DPOT-FT represents training for the same number of steps but using pre-trained weights for initialization. We have the following observations from the results.
@@ 4.6 Ablation Experiments
Here, we conduct ablation studies by training small-scale models on PDEBench datasets and compare the averaged zero-shot performance on corresponding test datasets.
@@
Influence of noise injection level. Here, we discuss the impact of the magnitude of noise injection on network training and generalization. As mentioned before, we observed that introducing noise into the input during auto-regressive training enhances the model’s robustness and stability for long trajectories in the test set. We experiment with noise levels $\varepsilon$ from $55e-$ 5 to $25e-$ 2 and the results are shown in Table 4. We find that the training error increases with the intensity of the injected noise, indicating that data with added noise becomes more challenging to learn. However, the test error shows a tendency to decrease from $0.0178$ to $0.0152$ before increasing with the addition of noise. We could view the noise injection as a regularization, and an appropriate level of noise benefits the generalization.
@@ 5 Conclusion
In this paper, we proposed the DPOT for large-scale PDE pre-training. We designed the new auto-regressive denoising strategy and a new model architecture based on Fourier attention. It was scaled up to 1B on 10+ PDE datasets. Then, we conducted extensive experiments on many PDE benchmarks and downstream PDE tasks. Our work is a pioneering work in large-scale exploration of the scaling behavior of PDE pre-training. It shows great potential for future applications as downstream tasks in practical problems.
@@ A.1 Model sizes and training details
Pre-training. We selected four models of varying sizes, i.e., DPOT-Tiny, DPOT-Small, DPOT-Medium, and DPOT-Large. The specific parameters of these models are shown in the table below. For the pre-training stage, we set the learning rate to $1\times 10^{-3}$ and used a One-cycle learning rate schedule over 1000 epochs, with the first 200 epochs as the warm-up phase. The AdamW optimizer was employed with a weight decay of $1\times 10^{-6}$ and momentum parameters $(\beta_{1},\beta_{2})=(0.9,0.9)$. Training was conducted using eight A800 GPUs, with a total batch size of 160. The patch size was set to 8. For datasets with long trajectories, such as PDB-DR and PDB-SWE, we assigned a weight of $w=3$, while for others, the weight was set to 1. For our training process, we selected $T=10$ timesteps to predict the next frame, aligning with the original settings of most datasets.
@@ Appendix B Supplementary Experimental Results
B.1 Results of DPOT with 1B Parameters
We have introduced a model with 1 billion parameters (DPOT-Huge), making it the largest among all current PDE pre-training models and an order of magnitude larger than all other models. We tested its zero-shot performance, and the results, compared with MPP-Large and DPOT-L, are shown in the Table 6. We observed significant improvements with DPOT-H as the model’s parameters increased, with performance boosts of over 50% on some datasets (like FNO datasets, PDEArena-NS) compared to DPOT-L. Except for a slight decrease in performance on the CFDBench dataset, DPOT-H nearly dominates the results.
L2RE MPP-L(0.4B) DPOT-L (0.5B) DPOT-H(1B)
FNO-1e-5 0.0550 0.0174
FNO-1e-4 0.0274 0.01314
FNO-1e-3 0.00528 0.00229
PDB-(1, 0.1) 0.0100 0.00961
PDB-(1, 0.01) 0.0216 0.0180
PDB-M1 0.0208 0.0158 0.0138
PDB-(0.1, 0.1) 0.00872 0.00847
PDB-(0.1, 0.01) 0.0115 0.0105
PDB-M0.1 0.0147 0.0101 0.00948
PDB-DR 0.0098 0.0232 0.0191
PDB-SWE 0.00220 0.00233 0.00199
PDEArena-NS 0.0798 0.0379
PDEArena-NS-cond 0.240 0.213
CFDBench 0.00650 0.00749
Table 6: Zero-shot performance comparison for 1B DPOT model.
B.2 Timing experiments
@@
B.3 Comparison with AFNO architecture
@@
B.4 Performance under varying resolutions
@@
B.5 Supplementary Experiments on Datasets with Long Trajectories
We added a supplementary experiment on long trajectory problems using DPOT. We generated 1000 data with a trajectory length of 500 at a 128x128 resolution, using the FNO-1e-3 dataset generation code. The model was trained and tested on subsets of 20, 50, 100, 200, and 500, recording the corresponding L2RE as shown in Table 10. We observed that at very short step sizes, the improvement brought by pre-training is minimal due to already high prediction accuracy. However, as the trajectory length increases, the prediction difficulty significantly rises. The pre-trained DPOT’s prediction error increases very slowly over time, whereas the error for the randomly initialized model grows quickly. This suggests that the pre-trained DPOT model can better utilize the prior knowledge of other PDEs to simulate the evolution of fluid over long trajectories.
Num steps 20 50 100 200 500
DPOT 0.00188 0.00192 0.00592 0.0241 0.0912
DPOT-pretrained 0.00148 0.00186 0.00335 0.0110 0.0385
Table 10: L2RE for DPOT and DPOT-pretrained at different trajectory steps
B.6 Additional Results on SuperBench Dataset
We conducted additional experiments on the SuperBench dataset, specifically on the Navier-Stokes Kraichnan Turbulence Flow problem (Re=16000), with the task of upsampling data by 8x, i.e., from 128x128 to 1024x1024. Following the SuperBench paper, we used RFNE (lower score means better performance) as the metric for comparison. The results are as shown in the Table 11, with SwinIR, WDSR, and EDSR results coming directly from the SuperBench original paper. We found that the non-pretrained DPOT network performs comparably to the current best model, EDSR, while the pre-trained DPOT model significantly outperforms EDSR. This demonstrates that pre-trained PDE models can enhance performance on downstream super-resolution tasks.
Model RFNE (%)
SwinIR 0.80
WDSR 0.72
EDSR 0.57
DPOT-vanilla 0.59
DPOT-pretrained 0.46
Table 11: RFNE (%) on the SuperBench dataset for the Navier-Stokes Kraichnan Turbulence (Re=16000) x8 Upsampling Task
B.7 Supplementary experiments on Kolmogorov turbulence flow
Additionally, we also added a dataset for Kolmogorov turbulence flow, a dataset with notable chaotic behavior. We generated 128 training trajectories with a resolution of 256x256x200 for training and 16 for testing. The results are shown in Table 12. We found that the vanilla DPOT model and the FNO model have a full trajectory test error of 82.2% and 104%, indicating that the model almost learned nothing, especially as time progressed. In contrast, the DPOT-pretrained error was 33.5%, significantly better than the non-pretrained DPOT model, suggesting it has learned some patterns of chaotic fluid evolution. It shows that pre-trained initialization is very helpful for complex, chaotic fluid simulations.
L2RE Train loss (step-wise) Test loss (total trajectory)
FNO 0.0931 1.04
DPOT 0.0485 0.822
DPOT-pretrained 0.0296 0.335
Table 12: Test results for Kolmogorov turbulence flow
B.8 Comparison between Equal Data Sampling and Our Balanced Data Sampling
We evaluated the impact of uniform sampling from all datasets combined versus our balanced sampling approach for pre-training. We retrained DPOT-Tiny ($\sim$7M) for comparison with the original model. The results are shown in Table 13. Our sampling method resulted in more even convergence across different datasets. For instance, with equal sampling, some datasets showed very poor convergence, such as FNO-1e-5, PDB-DR, PDB-SWE, etc. Although the performance on a few datasets like PDB-M1 was slightly better with equal sampling, overall, balanced sampling outperformed equal sampling.
DPOT-Tiny(L2RE) Equal sampling Balanced sampling
FNO-1e-5 0.177 0.0976
FNO-1e-4 0.0701 0.0606
FNO-1e-3 0.0327 0.00954
PDB-(1, 0.1) 0.0149 0.0173
PDB-(1, 0.01) 0.0376 0.0397
PDB-M1 0.0265 0.0285
PDB-(0.1, 0.1) 0.0135 0.0132
PDB-(0.1, 0.01) 0.0236 0.0220
PDB-M0.1 0.0185 0.0176
PDB-DR 0.170 0.0321
PDB-SWE 0.0225 0.00560
PDEArena-NS 0.126 0.125
PDEArena-NS-cond 0.407 0.384
CFDBench 0.00981 0.00952
Table 13: Performance comparison of DPOT-Tiny with Equal and Balanced Sampling.
B.9 Comparison between Self-attention and Fourier Mixer
We compared transformers using standard self-attention as a mixer with our DPOT model using a Fourier mixer, with results shown in Table 14. The datasets used are consistent with the paper’s ablation study, combining subsets of the Compressible Navier-Stokes equations from PDEBench. We found that models with self-attention require more parameters and computational resources. However, they do not fit or generalize to PDE data as well as models based on the Fourier mixer. Compared to self-attention mixers, the Fourier mixer could better utilize the spectrum properties of PDE data and reduce test errors by about 25%.
L2RE #N params Flops per forward Train loss Test loss
Self attention 52.1M 88.45G 0.0183 0.0238
Fourier mixer 30.8M 75.44G 0.0127 0.0174
Table 14: Performance comparison of Self-attention and Fourier mixer in DPOT model.
@@ Appendix C Proof of Universal Approximation of Fourier Attention
In the following section, we will demonstrate the Universal Approximation capability of DPOT. This means that the model possesses ample expressive ability, capable of fitting a large class of operators and adapting to the required precision. Specifically, we show the core layers, i.e. Fourier attention layers have the capacity to approximate any continuous operators.
@@ C.1 Preliminaries and Definition
Firstly, as DPOT mainly focuses on PDE problems defined on a rectangular domain, we will concentrate on functions defined on periodic torus $\mathbb{T}^{d}=[0,2\pi]^{d}$.
Given the domain, we can define Fourier transform for any function $v\in L^{2}(\mathbb{T}^{d})$ as:
$\mathcal{F}(v)(\xi)=\frac{1}{(2\pi)^{d}}\int_{\mathbb{T}^{d}}v(x)e^{-i\langle\xi,x\rangle}\text{d}x$ (22)
However, in the context of numerical calculation, we have to represent the function in a discrete way. In DPOT, we choose a regular mesh :=$\mathcal{M}:=\{x_{\mathbf{j}}\}\subset\mathbb{T}^{d}$, where the index vector is denoted as $\mathbf{j}=(j_{1},\cdots,j_{d})\in\mathcal{I}$ with =$\mathcal{I}=\{(j_{1},\cdots,j_{d})|\forall l\leq d,j_{l}\leq n\}$. Given the mesh, we can represent the function by its value on the mesh: $\{v_{\mathbf{j}}=v(x_{\mathbf{j}})\}$ and then define Discrete Fourier transform as:
($\mathcal{F}(\{v_{\mathbf{j}}\})(\mathbf{k})=\frac{1}{\sqrt{n^{d}}}\sum_{j_{l}\leq n}v_{\mathbf{j}}e^{-\frac{2\pi i\langle\mathbf{j},\mathbf{k}\rangle}{n}},\forall\ \mathbf{k}\in\mathcal{I}$ (23)
Similarly, we can define Discrete Inverse Fourier Transform as:
($\mathcal{F}(\{v_{\mathbf{j}}\})(\mathbf{k})=\frac{1}{\sqrt{n^{d}}}\sum_{j_{l}\leq n}v_{\mathbf{j}}e^{\frac{2\pi i\langle\mathbf{j},\mathbf{k}\rangle}{n}},\forall\ \mathbf{k}\in\mathcal{I}$ (24)
With the preparations in place, we can now formally define our model DPOT mathematically.
Definition C.1 (DPOT).
@@
A DPOT is a neural operator $\mathcal{N}$ defined between input space $\mathcal{X}^{n}$ and output space $\mathcal{Y}^{n}$, with $\mathcal{X}\subset\mathbb{R}^{d_{i}},\mathcal{Y}\subset\mathbb{R}^{d_{o}}$ being the range of input and output function. The structure of DPOT with $L$ layers can be described as:
$\mathcal{N}=\mathcal{Q}\circ\mathcal{L}_{L}\circ\cdots\circ\mathcal{L}_{1}\circ\mathcal{R}$ (25)
@@
where $L_{l}$ is the $l$-th fourier attention layers of DPOT, $\mathcal{R}$ is the lifting operator and $\mathcal{Q}$ is the projection operator, which will be defined below.
Definition C.2 (fourier attention layers of DPOT).
The $l$-th fourier attention layers of DPOT $L_{l}:\mathbb{R}^{d_{h}}\to\mathbb{R}^{d_{h}}$ is a mapping between hidden space with the structure:
$\mathcal{L}_{l}(\{v_{\mathbf{j}}\})(\mathbf{k})=v_{\mathbf{k}}+M_{l}\circ\mathcal{F}^{-1}\circ K_{l}\circ\mathcal{F}(\{v_{\mathbf{j}}\})$ (26)
Here, $M_{l}:\mathbb{R}^{d_{h}}\to\mathbb{R}^{d_{h}}$ is a mapping in hidden space, which is a 2-layer MLP with GELU activation $\sigma$. $K_{l}:\mathbb{C}^{d_{h}}\to\mathbb{C}^{d_{H}}$ is a mapping in the frequency domain of hidden space, which is also a 2-layer MLP with GELU activation $\sigma$.
Definition C.3 (lifting operator of DPOT).
The lifting operator of DPOT $\mathcal{R}$ is a mapping from the codomain of input function $\mathcal{X}$ into hidden space $\mathcal{R}^{d_{h}}$
($\mathcal{R}(\{v_{\mathbf{j}}\})(\mathbf{k})=R(v_{\mathbf{k}},\mathbf{k}),R\in C_{\infty}(\mathcal{X}\oplus\mathcal{I},\mathbb{R}^{d_{h}}),\forall\ \mathbf{k}\in\mathcal{I}$ (27)
where $R$ is the pointwise lifting function and $P$ is the positional encoding.
Definition C.4 (projection operator of DPOT).
the projection operator of DPOT $\mathcal{Q}$ is a mapping from hidden space $\mathbb{R}^{d_{h}}$ into the codomain of output function $\mathcal{Y}$:
$\mathcal{Q}(\{v_{\mathbf{j}}\})(\mathbf{k})=Q(v_{\mathbf{k}}),Q\in C_{\infty}(\mathbb{R}^{d_{h}},\mathcal{Y}),\forall\ {\mathbf{k}}\in\mathcal{I}$ (28)
where $Q$ is the pointwise projection function.
Then, we will introduce some key definitions and theorems used in the following proof.
Definition C.5 (Equivariant Function).
(Alberti et al., 2023)
For any sequence-to-sequence function $f:\mathcal{X}^{n}\to\mathcal{Y}^{n}$, with $\mathcal{X},\mathcal{Y}\subset\mathbb{R}^{d}$, is called an equivariant function, only when it is equivariant to the input order. That is, for any permutation $\pi:[n]\to[n]$, the function satisfies:
$f([x_{\pi(1)},\cdots,x_{\pi(n)}])=[f_{\pi(1)}(X),\cdots,f_{\pi(n)}(X)]$ (29)
Definition C.6 (Sumformer).
(Alberti et al., 2023) Let $\mathcal{X},\mathcal{Y}\subset\mathbb{R}^{d}$ and $d^{\prime}\in\mathcal{N}$, and let two functions $\phi:\mathcal{X}\to\mathbb{R}^{d^{\prime}},\psi:\mathcal{X}\times\mathbb{R}^{d^{\prime}}\to\mathcal{Y}$. A Sumformer is a function $S:\mathcal{X}^{n}\to\mathcal{Y}^{n}$ which can be evaluated by:
$S([x_{1},\cdots,x_{n}]):=[\psi(x_{1},\Sigma),\cdots,\psi(x_{n},\Sigma)]$ (30)
where,
$\Sigma:=\sum_{k=1}^{n}\phi(x_{k})$ (31)
In (Alberti et al., 2023), the authors demonstrates the following theorem:
Theorem C.7 (Universal Approximation by Sumformer).
For any Equivariant function $f:\mathcal{X}^{n}\to\mathcal{Y}^{n}$ and any $\varepsilon>0$, there exists a Sumformer $S$ that satisfies:
$\sup_{X\in\mathcal{X}^{n}}\|f(X)-S(X)\|_{\infty}<\varepsilon$ (32)